A garden table and a bench cost 855 combined. The garden table costs 45 less than the bench. What is the cost of the bench?
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52894)

.
Table + Bench = 855

Table = Bench - 45


(Bench - 45) + Bench = 855


2*Bench = 855 + 45 = 900


  Bench            = 900/2 = 450.     ANSWER

Solved.

I formatted my solution in a way it would be clear and self-explained without using words.



Answer by josgarithmetic(39630)